OpenStudy (anonymous):

Help please! Everything is in degrees so Sin 57 cos 13 - cos 57 sin 13

hero (hero):

\[\sin 57 \cos 13 - \cos 57 \sin 13 = \sin(57 - 13) = \sin(44)\]